6.5 kg chunk of ice at -120C is placed in a 0.22 kg of water at 220C. Ignore any heat flow with the surroundings, including the container, assume the final mixture is ice. Determine the following; 

heat lost or gained by water    


1
Expert's answer
2021-11-21T17:31:51-0500

"c_{ice}m_{ice}(t-t_{ice})=c_{water}m_{water}(t_1-t_2)+\lambda m_{water}+c_{ice}m_{water}(t-t_{ice})"


"2140\cdot 6.5\cdot (t-(-12))=4200\cdot 0.22\cdot(22-0)+333500\cdot 0.22+2140\cdot 0.22\cdot(0-t)"


So, we get "t=-5.1\u00b0C" .


"Q=4200\cdot 0.22\cdot(22-0)+333500\cdot 0.22+2140\cdot 0.22\cdot(0-(-5.1))="


"=96099.08\ (J)" . Answer